Two-point charges, q1, which has a charge of -76.11 uC and q2, which has a charge of 13.33 uC, are separated by a distance of 24.67 cm. Identify the field's magnitude at point A between the two charges if A is 10.39 cm away from the negative charge.


1
Expert's answer
2022-04-06T14:03:34-0400

We know that


"E_1=\frac{kq_1}{r^2}\\E_1=-\frac{9\times10^9\times76.11\times10^{-6}}{0.1039^2}=-634.53\times10^{5}N/C"

Electric field due to charge q2


"E_2=\frac{kq_2}{r^2}\\E_2=\frac{9\times10^9\times13.33\times10^{-6}}{0.3506^2}=9.759\times10^5N/C"

"E_{net}=E_1+E_2"


"E_{net}=(-634.53+9.759)\times10^5=-624.77\times10^{5}N/C"